OpenStudy (anonymous):

Which equation is in slope-intercept form and represents a line with slope through the point (0, 0)?

OpenStudy (akashdeepdeb):

I'll just tell you and then you can find it out yourself. y = mx + b YES y= mx +b THAT IS THE SLOPE INTERCEPT FORM A OF A LINE! m gives the lines slope and b gives the lines y-intercept! So if the line passes through the origin [YOUR CASE] Then it has a y-intercept (b) of 0 So the line should probably be like this y = mx [Because of 0 y -intercept] Understood? :) I hope you can get your answer now!! :D

OpenStudy (austinl):

We have the form \(y=mx+b\) We are not given a slope so we shall just leave in the variable "m". Next we will plug in the points in the coordinates given. \((x,y)\rightarrow(0,0)\) \(0=m(0) + b\) \(0=0+b\) \(b=0\) \(y=mx+0\rightarrow y=mx\)

OpenStudy (anonymous):

m always is that ! Y=m.x